A 100-kg body moving at 35 kph was found to have a total energy of 570 J. How much potential energy was there in the moving body?

a. 84 J


b. 490 J


c. 1.1 kJ


d. 55 J


1
Expert's answer
2021-05-26T07:13:21-0400

The energy is conserved, which means that the sum of a single body's kinetic and potential energies remains constant. Thus, we calculate the potential energy using the equation for the kinetic energy and the proper conversions (since we have the velocity in kph):


"E_T = E_{potential} + E_{kinetic}"


"E_{potential} = E_T - E_{kinetic} \, = E_{kinetic} - \frac{1}{2}mv^2 \,"


"E_{potential} = 570\,J - (\frac{100\,kg}{2})[(35\frac{km}{h})(\frac{1\,h}{3600\,s})(\frac{1000\,m}{1\,km})]^2"


"E_{potential} = [570 - (50)(\frac{350}{36})^2]\,J = [570 - 4726]\,J = -4156\,J"


I didn't reach an answer that matches the options given, I found the potential energy for the moving body to be Epotential = - 4156 J.


References

  • Ed. by J. J. Lagowski. (2004) Chemistry: Foundations and Applications. New York: Macmillan.